# NS-Net: Decoupling CLIP Semantic Information through NULL-Space for Generalizable AI-Generated Image Detection

### v1

@@ Abstract
The rapid progress of generative models, such as GANs and diffusion models, has facilitated the creation of highly realistic images, raising growing concerns over their misuse in security-sensitive domains. While existing detectors perform well under known generative settings, they often fail to generalize to unknown generative models, especially when semantic content between real and fake images is closely aligned. In this paper, we revisit the use of CLIP features for AI-generated image detection and uncover a critical limitation: the high-level semantic information embedded in CLIP’s visual features hinders effective discrimination. To address this, we propose NS-Net, a novel detection framework that leverages NULL-Space projection to decouple semantic information from CLIP’s visual features, followed by contrastive learning to capture intrinsic distributional differences between real and generated images. Furthermore, we design a Patch Selection strategy to preserve fine-grained artifacts by mitigating semantic bias caused by global image structures. Extensive experiments on an open-world benchmark comprising images generated by 40 diverse generative models show that NS-Net outperforms existing state-of-the-art methods, achieving a 7.4% improvement in detection accuracy, thereby demonstrating strong generalization across both GAN- and diffusion-based image generation techniques.
Introduction
@@
The rapid advancement of artificial intelligence has significantly accelerated the development of generative models, such as GANs (Karras et al. 2018; Park et al. 2019; Huang et al. 2024) and diffusion models (Ho, Jain, and Abbeel 2020; Rombach et al. 2022; Wu et al. 2024b). While these models are widely employed to generate highly realistic images and have enabled numerous beneficial applications, their potential misuse (Li et al. 2025c) raises serious concerns related to economic stability, political manipulation, and social security. Consequently, the reliable detection of AI-generated images has emerged as a critical research focus. To date, several effective detection methods (Zhao et al. 2021; Wu et al. 2024a; Tan et al. 2024a; Tao et al. 2025; Li et al. 2025a) have been proposed, particularly in scenarios where the generative model is known. However, developing robust and generalizable detection techniques that can effectively identify images from unknown generative models remains a fundamental and unresolved challenge.
Recently, a number of studies on AI-generated image detection (Ojha, Li, and Lee 2023; Liu et al. 2024a; Fu et al. 2025) have sought to enhance generalization by directly leveraging or fine-tuning CLIP (Radford et al. 2021)—a multimodal model pre-trained on large-scale image-text datasets, known for its strong generalization capabilities across unseen visual and textual data. Although the effectiveness of CLIP in this context has been demonstrated through quantitative experiments, there remains a lack of in-depth investigation into the intrinsic properties of its extracted features and how these characteristics influence detection performance.
Inspired by VIB-Net (Zhang et al. 2025), which aims to decouple category-related features, we revisit the distinctions between features with different intrinsic properties extracted by CLIP. Specifically, we construct two datasets, each containing 1,000 real images from LSUN and 1,000 images generated by SDXL. In the matched dataset, the generated images are conditioned on BLIP-derived textual descriptions of the corresponding real images. In contrast, the mismatched dataset consists of randomly selected real images and generated images conditioned on unrelated text prompts. We extract visual features using CLIP’s image encoder and visualize them using t-SNE, as shown in Figure 1(a). The features from the mismatched dataset exhibit clear clustering, suggestive of a linearly separable binary classification problem. In contrast, features from the matched dataset are more entangled and less distinguishable. This suggests that CLIP effectively separates real and generated images when semantic alignment is weak but struggles when the semantic content of both is closely aligned. Figure 1(b) presents the corresponding visualization of semantic features extracted by CLIP’s text encoder. Based on these observations, we propose the following hypothesis: The semantic information embedded in the visual features extracted by CLIP negatively impacts the performance of AI-generated image detection.
@@
To decouple the semantic information embedded in visual features, we construct the NULL-Space (Fang et al. 2024) corresponding to the semantic components in CLIP’s image features. The NULL-Space, formally defined as the set of vectors mapped to the zero vector under a given linear transformation, enables the removal of specific feature components. Given that CLIP is trained for text-image alignment, we leverage text features as a more tractable and semantically explicit representation, rather than attempting to disentangle the complex semantics directly from image features. Specifically, we extract features from the associated text prompts and compute their NULL-Space. Visual features extracted by CLIP’s image encoder are then projected onto this NULL-Space, thereby eliminating semantically aligned information and retaining components more relevant to forgery detection. As illustrated in Figure 1(c), this projection—achieved without any additional training—results in a clear separation between real and generated images.hile reducing the influence of semantic content on forgery detection (Zheng et al. 2024).
@@
To this end, we propose an innovative method, termed NS-Net, designed to extract more generalizable artifact information. Specifically, we leverage NULL-Space projection to decouple high-level semantic information encoded in CLIP features, followed by contrastive learning to model the underlying distributional differences between real and generated images. This strategy improves generalization ability beyond simple classification accuracy. Moreover, traditional center-cropping operations (Li et al. 2025b) often discard critical forgery traces—especially when manipulations occur away from the image center. To mitigate this, we introduce a Patch Selection strategy. Each image is first divided into uniform patches, from which those with the highest and lowest texture richness (based on predefined metrics) are selected. These selected patches are then shuffled and rearranged to form a new input image of the desired size. This process reduces the dominance of semantic content and enhances the preservation of fine-grained forgery cues. As shown in Figure 1(d), our method effectively removes task-irrelevant semantic features while retaining the generator-specific artifact patterns crucial for reliable detection. The main contributions of this work are summarized as follows:
@@
  • We innovatively construct the NULL-Space of the semantic features of generated images, which simply and efficiently eliminates task-irrelevant semantic information embedded in visual features extracted by image encoders, thus improving the generalization detection ability of AI-generated images.
  • We design a data preprocessing strategy called Patch Selection, which not only reduces the loss of forged information caused by center cropping operation adopted in previous work, but also utilizes patch rearrangement to further weaken the interference of irrelevant semantic information on detection performance.
  • Comprehensive experimental results have verified that our method exhibits strong generalization ability across 40 different generative models. Compared with the state-of-the-art methods, the detection accuracy is improved by 7.4%, which further highlights its superiority and universality.
@@
Related Work
Image Generation
With the rapid advancement of image generation technologies, the quality of synthetic images has significantly improved, enabling a wide range of practical applications. Generative Adversarial Networks (GANs) (Goodfellow et al. 2014) introduces adversarial training between a generator and discriminator, leading to substantial gains in image realism (Karras 2019; Park et al. 2019). More recently, diffusion models (Ho, Jain, and Abbeel 2020; Song, Meng, and Ermon 2020) have emerged as a powerful alternative, generating high-resolution and diverse images through a progressive denoising process (Ramesh et al. 2022; Rombach et al. 2022), while offering greater training stability than GANs. Beyond unconditional synthesis, techniques such as customized generation (Wu et al. 2024b; Ye et al. 2023) and image inpainting (Brooks, Holynski, and Efros 2023) have gained traction, enabling users to guide the generation process using textual or conditional inputs, which enhance content fidelity and better accommodate user-specific requirements.
AI-Generated Image Detection
To mitigate the societal risks associated with AI-generated images, researchers (McCloskey and Albright 2018; Farid 2022; Haliassos et al. 2021; Chen et al. 2022) have increasingly focused on developing effective detection methods. As generative models have advanced, recent researches have shifted toward improving the generalization capabilities of detection models. Some methods extract subtle artifacts from fine-grained spatial features, utilizing limited receptive fields (Chai et al. 2020; Cavia et al. 2024), modeling local pixel relationships (Tan et al. 2024b), or employing patch-based analysis (Chen and Yang 2020; Baraldi et al. 2025). Others explore frequency-domain artifacts to enhance generalization across generative models (Luo et al. 2021; Tan et al. 2024a; Yan et al. 2025). In addition, LGrad (Tan et al. 2023) leverages gradient signals from pre-trained models, while SAFE (Li et al. 2025b) combines cropping with data augmentation techniques such as ColorJitter, RandomRotation, and RandomMask to improve detection performance against sophisticated generation methods.
In recent years, many researchers have explored leveraging CLIP (Radford et al. 2021) for detecting AI-generated images. UnivFD (Ojha, Li, and Lee 2023) utilize CLIP’s image-text features for binary classification, demonstrating promising generalization capabilities. However, subsequent studies (Liu et al. 2024a, b; Fu et al. 2025) reveal that CLIP’s native features are not fully optimized for forgery detection, prompting efforts to fine-tune the model to better capture subtle artifacts. To this end, Fatformer (Liu et al. 2024a) introduces a forgery-aware adapter that integrates both spatial and frequency-domain extractors. C2P_CLIP (Tan et al. 2025) further enhances CLIP’s detection capability by injecting category-consistent cues into the text encoder. Additionally, VIB-Net (Zhang et al. 2025) addresses the issue of category interference in CLIP’s features by employing a Variational Information Bottleneck framework.
Despite ongoing advances, the intrinsic properties of CLIP features remains insufficiently understood. We hypothesize that the entanglement of semantic information within these features impairs the model’s ability to capture task-specific artifact cues essential for forgery detection.
Methodology
@@
We propose the NULL-Space Network (NS-Net), a novel framework designed to decouple semantic information embedded in CLIP features, preserving only the components relevant to forgery detection, which significantly improves the model’s generalization ability. As illustrated in Figure 2, our architecture begins with a Patch Selection strategy tailored to accommodate the fixed input size of CLIP, ensuring that potential forgery traces are retained. We then project the visual features extracted from CLIP’s image encoder onto the NULL-Space of the semantic information, thereby effectively removing irrelevant or confounding semantics. The resulting task-specific features are subsequently leveraged in a contrastive learning setup, which not only guides the linear classification layer but also models the intrinsic distributional differences between real and AI-generated images, beyond simple categorical prediction. The following sections elaborate on the proposed methodology in detail.
Patch Selection
Inspired by SAFE (Li et al. 2025b), which highlights that the resizing operations commonly used during training and testing can suppress artifact information, we incorporate a cropping strategy into the image pre-processing pipeline. However, conventional center cropping may inadvertently discard crucial forgery traces—especially when manipulated regions lie away from the image center. To address these issues, we propose the Patch Selection strategy that preserves task-relevant regions while adapting images to the fixed input size required by the network.
@@
Specifically, we first divide the image into multiple non-overlapping patches using a fixed window size, denoted as $I=\{x_{1},x_{2},...,x_{n}\},x_{i}\in{\mathbb{R}}^{N\times N\times 3}$. In this work, we set the patch size to $N=32$. Each patch is then transformed into the frequency domain using the Fast Fourier Transform (FFT), resulting in $I_{f}=\{x_{1}^{f},x_{2}^{f},...,x_{n}^{f}\},x_{i}^{f}\in{\mathbb{R}}^{N\times N\times 3}$. We compute the spectral entropy of each patch to quantify its texture complexity and sort all patches based on this value in descending order. Given the network’s required input image size $M\times M$, we determine the number of patches to be selected as $2K=(\frac{M}{N})^{2}$. We then select the top $K$ patches with the highest spectral entropy and the bottom $K$ patches with the lowest spectral entropy. These patches are randomly shuffled and reconstructed into a new image $X_{\text{{reconstruct}}}\in{\mathbb{R}}^{M\times M\times 3}$ to match the input requirements. The proposed patch selection strategy not only ensures compatibility with networks requiring fixed-size inputs, but also preserves both the texture-rich and texture-poor regions, thereby reducing the loss of critical artifact information. Moreover, the shuffling process disrupts the original spatial semantics of the image, helping the model focus on subtle artifacts rather than high-level semantic content—thus enhancing its ability to detect forgeries.
NULL-Space Decoupling
[…]

### v2

@@ Abstract
The rapid progress of generative models, such as GANs and diffusion models, has facilitated the creation of highly realistic images, raising growing concerns over their misuse in security-sensitive domains. While existing detectors perform well under known generative settings, they often fail to generalize to unknown generative models, especially when semantic content between real and fake images is closely aligned. In this paper, we revisit the use of CLIP features for AI-generated image detection and uncover a critical limitation: the high-level semantic information embedded in CLIP’s visual features hinders effective discrimination. To address this, we propose the NS-Net, a novel detection framework that leverages NULL-Space projection to decouple semantic information from CLIP’s visual features, followed by contrastive learning to capture intrinsic distributional differences between real and generated images. Furthermore, we design a Patch Selection strategy to preserve fine-grained artifacts by selecting patches with the highest and lowest spectral entropy. Extensive experiments on an open-world benchmark comprising images generated by 40 diverse generative models show that NS-Net outperforms existing state-of-the-art methods, achieving a 7.4% improvement in detection accuracy, thereby demonstrating strong generalization across both GAN- and diffusion-based image generation techniques.
[…]
1 Introduction
@@
The rapid advancement of artificial intelligence has significantly accelerated the development of generative models, such as GANs [karras2018progressive, park2019semantic, huang2024gan] and diffusion models [ho2020denoising, rombach2022high, wu2024infinite]. While these models are widely employed to generate highly realistic images and have enabled numerous beneficial applications, their potential misuse [li2025rethinking] raises serious concerns related to economic stability, political manipulation, and social security. Consequently, the reliable detection of AI-generated images has emerged as a critical research focus. To date, several effective detection methods [zhao2021multi, wu2024local, tan2024frequency, tao2025sagnet, li2025pay] have been proposed . However, developing generalizable detection techniques that can effectively identify images from unknown generative models remains a fundamental and unresolved challenge.
@@
Recently, a number of studies on AI-generated image detection [ojha2023towards, liu2024forgery, fu2025exploring] have sought to enhance generalization by directly leveraging or fine-tuning CLIP [radford2021learning]—a multimodal model pre-trained on large-scale image-text datasets, known for its strong generalization capabilities across unseen visual and textual data. Although the effectiveness of CLIP in this context has been demonstrated through quantitative experiments, there remains a lack of in-depth investigation into the intrinsic properties of its extracted features and how these characteristics influence detection performance.
Inspired by VIB-Net [zhang2025towards], which aims to decouple category-related features, we revisit the distinctions between features with different intrinsic properties extracted by CLIP. Specifically, we construct two datasets, each containing 1,000 real images from LSUN and 1,000 images generated by SDXL. In the matched dataset, the generated images are conditioned on BLIP-derived textual descriptions of the corresponding real images. In contrast, the mismatched dataset consists of randomly selected real images and generated images conditioned on unrelated text prompts. We extract visual features using CLIP’s image encoder and visualize them using t-SNE, as shown in Figure 1(a). The features from the mismatched dataset exhibit clear clustering, suggestive of a linearly separable binary classification problem. In contrast, features from the matched dataset are more entangled and less distinguishable. This suggests that CLIP effectively separates real and generated images when semantic alignment is weak but struggles when the semantic content of both is closely aligned. Figure 1(b) presents the corresponding visualization of semantic features extracted by CLIP’s text encoder. Based on these observations, we propose the following hypothesis: The semantic information embedded in the visual features extracted by CLIP negatively impacts the performance of AI-generated image detection.
@@
To decouple the semantic information embedded in visual features, we construct the NULL-Space [fang2024alphaedit] corresponding to the semantic components in CLIP’s image features. The NULL-Space, formally defined as the set of vectors mapped to the zero vector under a given linear transformation, enables the removal of specific feature components. Given that CLIP is trained for text-image alignment, we leverage text features as a more tractable and semantically explicit representation, rather than attempting to disentangle the complex semantics directly from visual features. Specifically, we extract features from the associated text prompts and compute their NULL-Space. Visual features extracted by CLIP’s image encoder are then projected onto this NULL-Space, thereby eliminating semantically aligned information and retaining components more relevant to forgery detection. As illustrated in Figure 1(c), without any additional training, this projection results in a clear separation between real and generated images while reducing the influence of semantic content on forgery detection.
@@
To this end, we propose an innovative method, termed NS-Net, designed to extract more generalizable artifact information. Specifically, we leverage NULL-Space projection to decouple high-level semantic information embedded in CLIP’s visual features, followed by contrastive learning to model the underlying distributional differences between real and generated images. This strategy improves generalization ability beyond simple classification accuracy. Moreover, the traditional CenterCrop [li2025improving], including PatchShuffle [zhong2023patchcraft, zheng2024breaking], only retains the artifacts at the center, overlooking critical forgery traces that may occur away from the center of the images. To mitigate this, we introduce a Patch Selection strategy. Each image is first divided into uniform patches, from which those with the highest and lowest entropy (as described in Sec.˜3.1) are selected. These selected patches are then rearranged to form a new input image of the desired size, which maximizes the preservation of artifact information. As shown in Figure 1(d), our method effectively removes task-irrelevant semantic features while retaining the generator-specific artifact patterns crucial for reliable detection. The main contributions are summarized as follows:
@@
  • We are the first to demonstrate that the semantic information embedded in the visual features extracted by CLIP adversely affects the performance of AI-generated image detection. To this end, we innovatively construct the NULL-Space of semantic features, which simply and efficiently eliminates task-irrelevant semantic information embedded in the image encoder’s visual representations, thus improving the generalization detection ability.
  • We design a data preprocessing strategy, Patch Selection, which maximizes the preservation of artifact information by selecting patches with the highest and lowest entropy.
  • Comprehensive experiments demonstrate that our method achieves superior generalization across 40 generative models, outperforming the state-of-the-art method by 7.4%.
2 Related Works
2.1 Image Generation
With the rapid advancement of image generation technology, the visual quality of synthesized images has improved dramatically, increasingly satisfying diverse application demands. Early Variational Autoencoders (VAEs) [kingma2013auto, van2017neural] generated images by learning the underlying data distribution; however, their outputs were often blurry and lacked fine details, which limits practical applicability. The introduction of Generative Adversarial Networks (GANs) [goodfellow2014generative] revolutionized the field by significantly enhancing image realism through adversarial training between generators and discriminators, which have been widely adopted in image synthesis [chen2020simswap] and style transfer [karras2019style, park2019semantic]. More recently, diffusion models [ho2020denoising, song2020denoising] have achieved remarkable success in generating high-resolution and highly diverse images through a progressive denoising process from noise to data, demonstrating strong potential in artistic and commercial applications [ramesh2022hierarchical, rombach2022high]. In addition, customized generation [wu2024infinite, ye2023ip] and image inpainting [brooks2023instructpix2pix] have emerged as new trends, allowing users to guide the generation process through textual or conditional inputs. Such controllable generation not only improves the accuracy and coherence of synthesized content but also caters to increasingly personalized user needs.
2.2 AI-Generated Image Detection
To mitigate the societal risks associated with AI-generated images, researchers have increasingly focused on developing effective detection methods. Early approaches primarily targeted semantic inconsistencies [mccloskey2018detecting, farid2022lighting, haliassos2021lips, chen2022self] in the spatial domain. However, as generative models have advanced, single-domain detection methods have become insufficient against the growing diversity of generative techniques. In response, recent research has shifted toward improving the generalization capabilities of detection models. Many of them focus on capturing generalizable low-level forgery clues, including local pixel relationships [chai2020makes, chen2020manipulated, tan2024rethinking, cavia2024real, baraldi2025contrasting], frequency-based features [luo2021generalizing, tan2024frequency, yan2025generalizable], gradients [tan2023learning], and reconstruction errors [wang2023dire, chen2024drct, yu2024semgir, chu2025fire]. To illustrate, F3-Net [qian2020thinking] introduces frequency component division and the frequency statistical difference between real and forgery images into face forgery detection; LGrad [tan2023learning] leverages gradient signals from pre-trained models; NPR [tan2024rethinking] reinterprets upsampling operations to construct an efficient artifact representation; DIRE [wang2023dire] extracts the reconstruction error of the image over a pre-trained diffusion model as the artifact of diffusion-based images. In addition, some studies adopt data-driven approaches to enhance model generalization, focusing on data preprocessing [wang2020cnn, chen2024drct, li2025improving] and the construction of augmented datasets [guillaro2025bias, yang2025d, chen2025dual], which improve the detection performance against sophisticated generation methods.
In recent years, many researchers have explored leveraging CLIP [radford2021learning] for detecting AI-generated images, owing to their powerful feature extraction and semantic understanding capabilities. UnivFD [ojha2023towards] utilizes CLIP’s image-text features for binary classification, demonstrating promising generalization capabilities. However, subsequent studies [liu2024forgery, liu2024mixture, fu2025exploring] reveal that CLIP’s native features are not fully optimized for forgery detection, prompting efforts to fine-tune the model to better capture subtle artifacts. To this end, Fatformer [liu2024forgery] introduces a forgery-aware adapter that integrates both spatial and frequency-domain extractors. C2P-CLIP [tan2025c2p] further enhances CLIP’s detection capability by injecting category-consistent cues into the text encoder. Additionally, VIB-Net [zhang2025towards] addresses the issue of category interference in CLIP’s features by employing a Variational Information Bottleneck framework; Effort [yan2024orthogonal] explicitly constructs two orthogonal subspaces for pre-trained knowledge and forgery features using singular value decomposition (SVD), achieving strong generalization performance.
Despite continuous advances, the intrinsic properties of CLIP’s visual features remain insufficiently understood. We hypothesize that the entanglement of semantic information within these features impairs the model’s ability to capture task-specific artifact cues that are essential for forgery detection, motivating our goal to decouple the embedded semantics in visual features.
2.3 NULL-Space
The NULL-Space [coleman1986null] refers to the set of vector subspaces that are mapped to the zero vector under a linear transformation, and has been widely applied in natural language processing (NLP) [ravfogel2020null, ravfogel2022linear]. In recent years, with the proliferation of large-scale pre-trained models, the concept of NULL-Space has also been increasingly adopted in continual learning [zeng2019continual, saha2021gradient, guo2022adaptive, fang2024alphaedit] to mitigate catastrophic forgetting. Specifically, OWM [zeng2019continual] projects new task updates onto the orthogonal complement of the past input subspace to prevent forgetting without replaying old samples, while AlphaEdit [fang2024alphaedit] projects parameter perturbations onto the NULL-Space of knowledge to be retained, substantially reducing interference and forgetting during sequential editing.
As discussed above, projecting new knowledge into the NULL-Space of old knowledge makes it orthogonal to the old knowledge, thus mitigating catastrophic forgetting. By analogy, We treat semantic features as a subspace whose influence on visual representations we aim to eliminate. Accordingly, we project visual features onto the NULL-Space of this semantic subspace, ensuring that visual and semantic components are orthogonal. This orthogonality implies that variations in semantic features do not affect the projected visual representations, evidencing effective decoupling. Different prior NULL-Space methods that preserve previously learned knowledge, we leverages the NULL-Space to explicitly eliminate task-irrelevant semantics, yielding more generalized and artifact-relevant representations.
@@ 2.3 NULL-Space
3 Methodology
@@
We propose the NULL-Space Network (NS-Net), a novel framework designed to decouple semantic information embedded in CLIP features, preserving only the components relevant to forgery detection, which significantly improves the model’s generalization ability. As illustrated in Figure 2, our architecture begins with a Patch Selection strategy to preserve fine-grained artifacts by selecting patches with the highest and lowest spectral entropy. We then project the visual features extracted from CLIP’s image encoder onto the NULL-Space of the semantic information, thereby effectively removing irrelevant or confounding semantics. The resulting task-specific features are subsequently leveraged in a contrastive learning setup, which not only guides the linear classification layer but also models the intrinsic distributional differences between real and AI-generated images, beyond simple categorical prediction. The following sections elaborate on the proposed methodology in detail.
3.1 Patch Selection
In practice, input images are often extremely high-resolution. Although Vision Transformers (ViTs) are theoretically size-agnostic, their quadratic complexity with respect to sequence length makes large-scale processing computationally prohibitive; convolutional networks face similar limitations. Many studies adopt Resize to downsample inputs to an appropriate resolution, but this process inevitably suppresses artifact information[li2025improving]. While CenterCrop [li2025improving, tan2025c2p, yan2024orthogonal] is commonly used to alleviate the information loss caused by resizing, it only retains the artifacts at the center, overlooking critical forgery traces that may occur away from the center of the images. Different generation mechanisms exhibit complementary spectral-entropy characteristics: GANs often introduce high-frequency periodic peaks and sharpening residues, whereas diffusion models produce low-frequency distortions such as over-smoothing. To address these issues, we propose the Patch Selection strategy that maximizes the preservation of complementary high/low-entropy artifacts while adapting images to the fixed input size.
@@ 3.1 Patch Selection
Specifically, we first divide the image into multiple non-overlapping patches using a fixed window size, denoted as $I=\{x_{1},x_{2},...,x_{n}\},x_{i}\in{\mathbb{R}}^{N\times N\times 3}$, where $n$ is the total number of patches. In this work, we set the patch size to $N=32$. Each patch is transformed into the frequency domain via Fast Fourier Transform (FFT), yielding $I_{f}=\{x_{1}^{f},x_{2}^{f},...,x_{n}^{f}\},x_{i}^{f}\in{\mathbb{R}}^{N\times N\times 3}$. We then compute the spectral entropy of each patch to quantify its texture complexity and sort all patches in descending order. Given the network’s required input image size $M\times M$, we determine the number of selected patches as $2K=(\frac{M}{N})^{2}$, where $n\geq 2K$. We then select the top $K$ patches with the highest spectral entropy and the bottom $K$ patches with the lowest spectral entropy from the original $n$ patches. These patches are randomly shuffled and reconstructed into a new image $X_{\text{{reconstruct}}}\in{\mathbb{R}}^{M\times M\times 3}$ to match the input requirements. The proposed patch selection strategy not only ensures compatibility with networks requiring fixed-size inputs, but also preserves both the texture-rich and texture-poor regions, thereby reducing the loss of critical artifact information. Moreover, the shuffling process disrupts the original spatial semantics of the image, helping the model focus on subtle artifacts rather than high-level semantic content, thus enhancing its ability to detect forgeries. Different from existing PatchShuffle preprocessing methods [zhong2023patchcraft, zheng2024breaking, gye2025reducing], which still use Crop or Resize before shuffle, we adopt a "selective alternative cropping" strategy that directly selects patches rich in artifact information, thereby reducing information loss at the source and further enhancing the model’s generalization ability.
3.2 NULL-Space Decoupling
Preliminarily, the NULL-Space of a given matrix $A$, denoted as $\text{NULL-Space}(A)$ , is defined as:
$\text{NULL-Space}(A)=\{X:AX=0\}.$ (1)